Acceleration due to gravity at the earth s surface is 9.8ms-2At what height above the earth is the g equal to 9.8/4ms-2


1
Expert's answer
2021-03-23T11:15:46-0400

The acceleration due to gravity at the Earth's surface can be written as follows:


"g_1=\dfrac{GM}{R^2}."

The acceleration due to gravity at the height "h" above the Earth can be written as follows:


"g_2=\dfrac{GM}{(R+h)^2}."

Let's divide "g_1"by "g_2":


"\dfrac{g_1}{g_2}=\dfrac{(R+h)^2}{R^2},""\sqrt{\dfrac{g_1}{g_2}}=\dfrac{(R+h)}{R},""h=R(\sqrt{\dfrac{g_1}{g_2}}-1),""h=6.37\cdot10^6\ m\cdot(\sqrt{\dfrac{9.8\ \dfrac{m}{s^2}}{2.45\ \dfrac{m}{s^2}}}-1)=6.37\cdot10^6\ m."